OpenStudy (anonymous):

Find the area and the circumference of a circle with radius 9cm.

OpenStudy (anonymous):

@Ryan_Kochan

OpenStudy (anonymous):

Circumference = 2 • π • radius = π • diameter Circle Area = π • r² = ¼ • π • d²

OpenStudy (anonymous):

\[A = \pi \times r^2\]

OpenStudy (anonymous):

\[C = 2 \times \pi \times r\]

OpenStudy (anonymous):

Circumference = 2 x 3.14 x 9 Area = 3.14 x 9^2

OpenStudy (anonymous):

A= 254.34 C= 56.52

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

would A be cm^2 and C be cm^3

OpenStudy (anonymous):

I think Not 100% sure on that one :P

OpenStudy (anonymous):

A : cm² and C: cm

OpenStudy (anonymous):

cm³ is used for volume